OpenStudy (beccab003):

How many ml of 10 percent acid should be added to pure acid to make 60 ml of 30 percent acid? Please explain answer.

hero (hero):

If: x = mL of 10% acid y = mL of pure acid Then: The Total mL of Acid is: x + y = 60 And: The Total Percent mL of Acid is: .10x + y = .30(60) So solve this system for \(x\) to find how much of the 10% acid you need to add: x + y = 60 .10x + y = 18

hero (hero):

Keep in mind that 100% = 100/100 = 1

OpenStudy (beccab003):

So what is the final result?

hero (hero):

You have to solve the system to find out.

OpenStudy (beccab003):

Will you tell me if I have the correct answer?

hero (hero):

It would be best for you to post your steps. You can use the draw button to write or you can just type them out.